### v2

# Explainability of Text Processing and Retrieval Methods: A Survey

Abstract.
Deep Learning and Machine Learning based models have become extremely popular in text processing and information retrieval. However, the non-linear structures present inside the networks make these models largely inscrutable. A significant body of research has focused on increasing the transparency of these models. This article provides a broad overview of research on the explainability and interpretability of natural language processing and information retrieval methods. More specifically, we survey approaches that have been applied to explain word embeddings, sequence modeling, attention modules, transformers, BERT, and document ranking. The concluding section suggests some possible directions for future research on this topic.
Explainability, Interpretability, Text Processing, Information Retrieval, Natural Language Processing, Machine Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrieval
1. Introduction
Given the volume of information that we routinely encounter, we have become acutely dependent on technology that processes the available information, classifies it on the basis of diverse criteria, selects useful information from this barrage, and presents it in a form that is easy to consume and assimilate. It is not enough for such technology to be merely effective: as humans, we would also like to understand why our tools behave the way they do. Naturally, a substantial body of research has focused on methods for providing high-level, human-understandable explanations or interpretations of the inner workings of, as well as the results produced by, information processing systems.
Up to about 10 years ago, commonly used information processing tools generally extracted a feature vector — a list of features along with their associated numerical weights — from the form in which the information was initially provided: structured or unstructured text, images, audio, video, etc. These features typically corresponded to human-understandable properties or attributes of the original information. We confine our discussion in this article to textual information only. For text, features would include words and phrases occurring in the text, their counts, morphology, parts of speech, and various grammatical relationships. These features were chosen on the basis of linguistic considerations; they were, in a sense, interpretable by design. A number of studies (some of which are discussed in more detail in Section 6) did focus on understanding feature weights, their behaviour, and the mathematical / statistical models underlying their computation and use, but they were not explicitly designated as explainability / intepretability studies.
Over the past few years, advances in Deep Machine Learning (ML) techniques have revolutionised the state of the art in textual information processing (as well as a number of other areas). These improvements have come at a cost: text is no longer simply represented in terms of easily interpretable features. Instead, early in the information processing pipeline, text is converted to so-called “dense” vectors in $\mathbb{R}^{n}$; the dimensions of this space are no longer easily understood. This representation is then processed by ML systems that have large and complex architectures, involving many parameters, often numbering in the millions, or even billions. As a fallout, there has been a marked increase in the attention devoted to providing high-level, intuitive explanations of these models and methods.
Thus, this seems to be an appropriate time to present an organised summary of the major research efforts in the area of explainability in textual information processing and retrieval. Indeed, several such general surveys have already been published (Søgaard, 2021; Madsen et al., 2022; Zini and Awad, 2022; Anand et al., 2022), in addition to more specialised surveys, such as (Belinkov and Glass, 2019; Dufter et al., 2022). Our objective is to complement these existing surveys by describing more recent work in areas already covered there, and by focussing on areas not addressed by them. Before discussing the relation between this survey and the existing ones in Section 2, we first briefly describe how we collected the set of papers reviewed in this survey.
1.1. Scope of this survey
Text processing and retrieval problems are studied by researchers from the Information Retrieval (IR) and Natural Language Processing (NLP) communities, as well as from the broader Machine Learning / Deep Learning (DL) communities. Thus, to ensure comprehensive coverage of explainability and interpretability studies relating to IR and NLP, we identified a wide range of highly-regarded venues including SIGIR, CIKM, WSDM, WWW, the ACL family of conferences and journals, AAAI, ICML, ICLR, NeurIPS, FAccT, ACM TOIS and IPM (Information Processing & Management).
The principal terms used in our search were variants of the keywords explanation (including explain*) and interpretability (viz., interpret*). These terms have been variously defined within the community. For example, according to (Ribeiro et al., 2016), explaining a model may be defined as presenting textual or visual cues that provide qualitative understanding of the relationship between the components of an input instance (e.g. words in text), and the model’s output for that instance. Doshi-Velez and Kim (2017) define interpretability more generally as simply “the ability to explain or to present in understandable terms to a human”. We adopt the broad sense conveyed by these definitions, and use the terms explainability and interpretability interchangeably in this survey.
Our preliminary search revealed a substantial body of work on explainability in Recommender Systems and Computer Vision. However, we excluded papers from these areas unless they directly contributed to follow-up work in IR or NLP. While IR systems and Recommender Systems share similarities in their goal of ranking items to meet users’ information needs, often with or without explicit queries, this survey focuses specifically on the retrieval of text documents. As text-focused IR methods are deeply intertwined with NLP techniques, we concentrate on the explainability of IR and NLP to maintain this focus. We include papers published from 2004 to 2024 in this set. Some influential works were also added that were published long before 2004. From the longlisted set of papers, a subset of 45 papers was selected for discussion in this article.
Organisation of the survey:  The remainder of this survey is organised as follows. Section 2 lists existing surveys that cover explainability in textual IR and NLP, and describes the relation between them and the current article. Section 3 introduces the basic terminology used in the literature on explainability, and presents a schema for classifying explainability research. The studies summarised in the rest of the article are broadly organised using this schema. Section 4 reviews some well-known approaches to explainability that are most commonly used as baselines in recent work. Section 6 is devoted to a detailed discussion of explainable document / passage ranking, and is itself divided into subsections by topic. Within each of these sections and subsections, different studies have been discussed in chronological order. Sections 7 through 10 are intended as supplements to existing surveys (specifically those by Madsen et al. (2022) and Zini and Awad (2022)); they cover recent work on explainability for individual elements of deep text processing pipelines, viz., word embeddings, sequence models, attention mechanisms, and transformer architectures. We conclude in Section 11 with a few research questions that we believe merit further investigation. 11todo: 1Update after everything is done.
2. Related Surveys
In a monograph (Lin et al., 2021), Lin et al. provide a comprehensive survey of neural IR models, covering topics from multi-stage ranking architectures to dense retrieval techniques. The authors observe that IR models may be stratified into three fairly well-separated levels on the basis of effectiveness: transformer-based ranking strategies represent a definite advance over pre-transformer Neural Ranking Models (NRMs), which in turn improve significantly over “traditional” [^1] In this survey, we use this adjective to refer to all non-neural IR models, including relatively recently proposed ones. statistical ranking models. An intuitive (but precise) grasp of what leads to these quantum jumps in effectiveness seems eminently desirable. This, along with a number of other factors, has led to much interest in the explainability of NRMs. There has always been a substantial overlap between the IR and NLP communities. With the recent success of neural methods, the boundary between the disciplines has been blurred even further. Not surprisingly, therefore, research in explainable IR (ExIR) is significantly influenced by similar research in NLP, which in turn draws heavily from explainability studies in ML in general.
Our search for a comprehensive summary of research on this broad subject led us to two recent articles that provide an overview of attempts to explain neural models for NLP (Madsen et al., 2022; Zini and Awad, 2022). In (Madsen et al., 2022), the main emphasis is on post-hoc [^2] In Section 3, we provide a review of this and other basic terms commonly used in the context of interpretability / explainability. methods for interpreting neural NLP methods, while Zini and Awad (2022) focus on the explainability of word embedding vectors, recurrent neural networks (RNNs) and attention mechanisms.
This being an area of active research, a non-trivial body of work exists that has not been covered by the above surveys, either because the work was published later, or because its topic was outside the specific scope of the surveys. In particular, until very recently, there were no similar overviews addressing ExIR in detail. The current study started as an attempt to address that lacuna. Recently, Anand et al. (2022) have published an article focused on this specific topic. There is, inevitably, a good deal of overlap between Anand et al.’s survey and Section 6 of this article, but we also cover a fair amount of work that is not discussed by them. Further, since our survey was prepared independently, we believe that it provides a somewhat different perspective on ExIR. Finally, given that ExIR often draws inspiration from explainability for NLP, we review (in Sections 7 through 10) recent work on Explainable NLP that was not covered by the earlier surveys (Madsen et al., 2022; Zini and Awad, 2022). In the rest of this section, we briefly discuss other related surveys that are either older, or have a scope that does not substantially overlap with ours.
To the best of our knowledge, Doshi-Velez and Kim (2017) were the first to attempt to formalise various notions related to interpretability in ML. Lipton (2018) provides an exposition of the broader aspects of interpretability in ML and Artificial Intelligence (AI), and why it is challenging, especially in the modern scenario. Søgaard (2021) examines extant taxonomies of explainability methods, and carefully constructs a framework for organising them. The next section draws from these articles to build a glossary of common terms used in this area.
Belinkov and Glass (2019) cover explainable NLP, and summarise studies that analyse neural network nodes, and activations across different layers of deep neural networks, in order to understand what linguistic features are encoded by neural NLP models. As mentioned above, many of the discussed methods were based on general interpretability approaches proposed by the ML community, often within the context of Computer Vision / Image Processing problems. Signalling the growing interest in this area, Danilevsky et al. (2020) also provide an early overview of explainability for NLP, describing different explanation categories and approaches, as well as several visualization techniques. Lertvittayakumjorn and Toni (2021) focus on explanation as a tool for debugging NLP models, with humans in the loop.
As Transformer and BERT-based techniques were widely adopted in the IR/NLP community, several researchers focused on various aspects of BERT, from its knowledge-representation power to fine-tuning, pre-training, and over-parametrisation issues. Rogers et al. (2020) were the first to survey research in this category, and present a summary of “what we know about how BERT works.” The surveys discussed at the beginning of this section provide a more up-to-date summary of research in this broad area.
There are other broader efforts in Explainable AI (XAI), such as XAIR (Clement et al., 2023), which systematically reviews XAI methods aligned with the software development process. However, this meta-review targets general AI explainability rather than the specific challenges of textual and retrieval systems addressed in our survey.
3. Categorising Explanation Methods
Methods for explaining IR/NLP models and techniques can be categorised on the basis of a number of criteria. While introducing these criteria (and also in the rest of this article), we use the symbols listed in Table 1.
$\mathcal{Q}$ set $\{{Q}_{1},{Q}_{2},\ldots,{Q}_{m}\}$ of queries
$\mathcal{D}$ corpus $\{{D}_{1},{D}_{2},\ldots,{D}_{n}\}$ of documents
$Q$ representative of any query from $\mathcal{Q}$, comprising a set ${q}_{1},{q}_{2},\ldots,{q}_{t}$ of terms
$D$ representative of any document from $\mathcal{D}$ consisting of the terms ${d}_{1},{d}_{2},\ldots,{d}_{l}$
$\mathcal{C}$ class of complex models that need explanations
$\mathcal{M}_{C}$ representative of any member of $\mathcal{C}$
$\mathcal{S}$ class of simple models relatively easily understood by humans
$\mathcal{M}_{S}$ representative of any instance of $\mathcal{S}$
II representative of any instance provided as input to a model
$I_{C}$ actual representation of $I$ (the “feature vector”, often hard to interpret) that is taken as input by $\mathcal{M}_{C}$
$I_{S}$ simple representation of $I$, easily understood by humans, often provided as input to $\mathcal{M}_{S}$
Table 1. Summary of notation used in this article.
In subsequent sections, we try to classify and label the reviewed approaches using the criteria described below.
3.1. Model-agnostic vs. model-aware explanations
Explanations that use information about the internal structure of a model $\mathcal{M}_{C}$ are referred to as model-specific or model-aware. In contrast, model-agnostic explanations treat the explained model as a black box, and make no assumptions about its internal architecture or parameters. Model-agnostic methods often submit controlled perturbations of an input instance $I$ to $\mathcal{M}_{C}$, and try to gather insights about $\mathcal{M}_{C}$ based on how its output is affected by these perturbations.
3.2. Inherent explainability vs. post-hoc interpretability
Some ML models, e.g., rule-based systems, decision trees and sparse linear models, are regarded as inherently explainable, because it is relatively simple for a human to trace and describe their inner workings. Traditional IR models like BM25 are also generally regarded as inherently explainable. Such models comprise the class $\mathcal{S}$ (see Table 1). Explanations of such models are, naturally, model-aware. The composition of $\mathcal{S}$ is open to debate, however. Moshkovitz et al. (2020) argue that even apparently well-understood methods with a single parameter, viz., $k$-means and $k$-medians clustering techniques, may be regarded as “opaque”, because these methods define cluster boundaries on the basis of linear combinations of features, rather than a single feature value. More pertinent to our context is the debate on whether the attention mechanism can be construed as providing explanations. Section 9.1 discusses this debate in more detail.
In contrast, post-hoc explanations primarily seek to elucidate the output, rather than the inner workings of models. Visualisations of word embeddings (e.g., using t-SNE (van der Maaten and Hinton, 2008)) are a form of post-hoc “explanation”: they do not explain why certain words are mapped to certain vectors, but once the vectors are generated (i.e., after the fact), they provide insights about the correspondence between proximity in the semantic and vector spaces. Explanations by example (e.g., “Model $M$ predicts class $Y$ for input instance $X$ because $X$ is ‘similar’ to training instances ${X}_{1},{X}_{2},\ldots,{X}_{k}$ which were labelled YY”) also fall within this category. Post-hoc explanations may be either model-aware or model-agnostic.
3.3. Scope of an explanation: global vs. local
Global methods explain the overall behavior of a model. Global explanations should apply to all (or most) input-output pairs for a given model. An example of a global explanation in the IR context is: “ranking based on query likelihood tends to be biased in favour of short documents when Jelinek-Mercer smoothing is used instead of Dirichlet smoothing” (Smucker and Allan, 2006). On the other hand, a local explanation only seeks to explain a model’s decision for a given input instance.
3.3.1. Pairwise vs. pointwise vs. listwise local explanations
For ranking models, local explanations can in turn be classified into three sub-categories. Pointwise explanations explain the specific decision made by a model for a particular input instance. In the IR context, a pointwise explanation addresses the question of why a document $D$ receives a particular similarity score for a query $Q$. Apart from ranking models, pointwise explanations may also be constructed for classifiers. Pairwise explanations are typically offered for systems that rank items (e.g., documents) based on their relationship to another item (e.g., a query). In the IR context, for a given query $Q$ and a document pair $\left<D_{i},D_{j}\right$>, a pairwise explanation tries to answer why a model ranks $D_{i}$ above/below $D_{j}$ for $Q$. Listwise explanations attempt to provide a single explanation for a complex model $\mathcal{M}_{C}$’s ranking of the top-$k$ items (documents) related to a given item (query). We discuss one possible form of such an explanation in the next section.
3.4. Nature or form of post-hoc explanations
At their simplest, explanations are observations or statements about how models behave. Some global explanations are of this form: see Section 3.3 for an example; more examples (about what knowledge BERT-based models possess) will be discussed in Section 6.5. Naturally, simple explanations can rarely be provided for complex neural models. In practice, therefore, the explanations constructed by researchers appear in diverse forms. Below, we briefly describe some forms in which explanations are commonly provided.
3.4.1. Explainable surrogate models
One widely adopted approach involves trying to construct an $\mathcal{M}_{S}\in\mathcal{S}$ that mimics the behaviour of a given complex model $\mathcal{M}_{C}$. In principle, $\mathcal{M}_{S}$, the explainable surrogate, may emulate $\mathcal{M}_{C}$ on most inputs; it may then be regarded as a global explanation. In practice, it is more feasible to provide a local explanation for a given $I$, by constructing an $\mathcal{M}_{S}\in\mathcal{S}$ along with a simple representation $I_{S}$ of $I$, such that the behaviour of $\mathcal{M}_{C}$ in the vicinity of $I$ is similar to the behaviour of $\mathcal{M}_{S}$ on $I_{S}$. As an example, suppose a neural ranker $\mathcal{M}_{C}$ generates a ranked list of documents for a given query $Q$. Then, a listwise explanation for this output might consist of a traditional document-ranking algorithm, e.g., BM25, and a list $L=\{{t}_{1},{t}_{2},\ldots,{t}_{k}\}$ of explanation terms. The pair $\left\langle\mathrm{BM25},L\right\rangle$ is interpreted as an explanation in the following sense: given $Q$, $\mathcal{M}_{C}$ must be implicitly detecting matches between documents and the terms in $L$ in the same way that BM25 would, if it were explicitly given $L$ instead of $Q$.
3.4.2. Feature attribution and rationales
Instead of explaining $\mathcal{M}_{C}$’s behaviour in terms of $\mathcal{M}_{S}$, researchers have also attempted to understand $\mathcal{M}_{C}$ by identifying which features of an input instance $I$ (i.e., which components of $I_{S}$ or $I_{C}$) most strongly affect $\mathcal{M}_{C}$’s output for $I$. This approach is called feature attribution. Several techniques for quantifying and visualising the contribution of a particular feature towards the prediction made by a model have been proposed in the literature. These local explanations may be provided graphically in the form of a heatmap or saliency map (Simonyan et al., 2014), with different colours or shades representing different degrees of importance of features, or simply as a list of the most important features along with their importance scores.
A related approach involves the construction of a rationale, a concise representation (or the extracted essence) of $I$ that is (i) both short and simple enough to be readily understandable by a human, but (ii) for which $\mathcal{M}_{C}$’s prediction is the same as that for $I$. For example, a short excerpt from an input text that is sufficient to correctly predict the sentiment of a sentence comprises a rationale for a sentiment classification model (DeYoung et al., 2020).
4. Basic Methods for Constructing Post-hoc Explanations
Before moving to a detailed discussion of recent explainability research in IR/NLP, we first review some well-established methods that are widely used as baselines against which more recent approaches are compared. These methods will be referred to repeatedly in the following sections and form a good starting point for exploring this sub-field.
4.1. Using probing classifiers to construct global explanations
Probing classifiers are being widely adopted for analyzing deep neural networks. They are based on the following idea. Let $\mathcal{M}_{C}$ be a neural model, trained for a specific target task, e.g., machine translation. Given a (preferably large, representative) set of inputs $\mathcal{I}=\{{I}_{1},{I}_{2},\ldots,{I}_{p}\}$, we collect the final (or an intermediate) representation of each $I_{j}$ — denoted by $I^{\prime}_{j}$ say, — generated within $\mathcal{M}_{C}$. A simple ‘probing’ classifier, $\mathcal{M}_{S}$, is trained to predict some specific linguistic property, e.g., part of speech (POS) of ${I}_{1},{I}_{2},\ldots,{I}_{p}$, using ${I^{\prime}}_{1},{I^{\prime}}_{2},\ldots,{I^{\prime}}_{p}$ as input. If $\mathcal{M}_{S}$ (the ‘probe’) performs well, we hypothesise that $\mathcal{M}_{C}$ has ‘learned’ the linguistic property in question by or before the layer from which the $I^{\prime}_{j}$s are extracted. If a probe uses only the representation generated by the penultimate layer, it may broadly be regarded as model-agnostic; any probe that provides insights into intermediate representations is, naturally, model-aware.
A typical probing framework would involve, for example, running a set of given sentences through a pre-trained version of BERT to generate contextual representations of the constituent tokens. A single-layer feed forward network could then be trained to take these embeddings as inputs, and predict the Part of Speech (POS) of each token. If the prediction accuracy of this POS tagger is high, we may claim that BERT’s representation encodes POS information. It should be clear from this discussion that probing is a model-aware method for providing global, post-hoc explanations.
While a number of studies (Köhn, 2015; Gupta et al., 2015; Hupkes et al., 2018) have used probing methods to evaluate hypotheses like “neural models learn syntactic structures”, Belinkov (2022) highlights the drawbacks of probing methods, some of which are described below.
  • There are no general guidelines for selecting appropriate baselines for probing experiments, and by implication, for determining whether the new classifier performs “well enough” to support a particular hypothesis.
  • Similarly, there is no well-defined process for determining whether the probing classifier is simple or complex, though some studies do suggest reporting the accuracy complexity trade-off, or deal with parameter free probing.
  • The effects of the size, composition, and relationship between the datasets used for the original and probing tasks have not been studied rigorously, and need to be carefully investigated.
  • Probing frameworks may be able to associate some linguistic property with an intermediate representation, but it does not clearly establish what role, if any, that particular property played in the prediction of the original classifier.
4.2. LIME: model-agnostic, local explanations
LIME (Ribeiro et al., 2016) is one of the earliest methods to be proposed in the current wave of research on explainability. Given an input $I$, LIME generates a set $\pi_{S}$ of slightly perturbed instances of $I_{S}$ (recall that $I_{S}$ denotes some interpretable representation of $I$: for a document, $I_{S}$ may be a binary term-vector), by randomly setting a small fraction of the nonzero elements of $I_{S}$ to zero. For each perturbed instance, the corresponding complex feature vector $I_{C}$ is also constructed; together, these comprise $\pi_{C}$. A simple classifier, e.g., a linear model, is trained on $\pi_{S}$ to mimic (and thus “explain”) $\mathcal{M}_{C}$ on $\pi_{C}$. No assumptions are made about $\mathcal{M}_{C}$. LIME is thus a model-agnostic (§3.1) method that provides simple surrogates (§3.4.1) as local (§3.3) explanations.
4.3. Feature attribution as model-agnostic, local explanations
As discussed in § 3.4.2, the objective of feature attribution is to quantify the contribution of a particular feature towards the output of $\mathcal{M}_{C}$ on $I$. Possibly the simplest feature attribution technique is occlusion (Zeiler and Fergus, 2014; Li et al., 2016) (sometimes referred to as the leave one out method). This estimates the importance of a group of features by setting all these features to zero, and measuring the resulting decrease in prediction accuracy. [^3] This general idea has been used in a wide variety of contexts where effects have to be measured. Since this approach only examines how the output of the model is affected by changes in the input, it is model-agnostic. Two other commonly-used families of feature attribution methods are discussed below. A library implementing many of these methods is available from https://captum.ai/, and may be used to interpret PyTorch models.
22todo: 2Possible to fit deeplift (deeplift) into one of the two following subsections?
4.3.1. Gradient based techniques
Sundararajan et al. (2017) proposed Integrated Gradients, a well-known method in this class of approaches. To quantify the importance of each feature, $I_{C}$ (the feature vector of an input instance $I$) is compared to a “baseline” ($I^{(0)}$, say) that corresponds to a very simple input. For example, if the input is an image, the baseline could be a black image; when studying word embeddings, the zero vector would be a possible baseline. We consider $m$ additional instances by linearly interpolating between $I_{C}$ and $I^{(0)}$. If $\mathbf{i}_{j}$ denotes the $j$-th feature or dimension of $I_{C}$, then its importance (or effect on the prediction) is measured in terms of its integrated gradient $IG_{j}(I_{C},I^{(0)})$, which depends on two factors:  (i) the difference between $\mathbf{i}_{j}$ and its corresponding value $\mathbf{i}^{(0)}_{j}$ in the baseline $I^{(0)}$,  and  (ii) $\frac{\partial f(I_{C})}{\partial\mathbf{i}_{j}}$, the gradient of the prediction $f$ with respect to $\mathbf{i}_{j}$. Instead of using the value of the gradient at a single point, we use the average of the gradients computed at the $m$ interpolation points between $I_{C}$ and $I^{(0)}$. More concretely,
$IG_{j}(I_{C},I^{(0)})=(\mathbf{i}_{j}-\mathbf{i}^{(0)}_{j})\times\frac{1}{m}\sum_{\alpha=1}^{m}\frac{\partial f(I^{(0)}+\frac{\alpha}{m}\times(I_{C}-I^{(0)}))}{\partial\mathbf{i}_{j}}.$
4.3.2. Shapley
The Shapley value (Shapley, 2016) was originally proposed in 1952 in the context of Cooperative Game Theory to quantify the contribution of a player to the outcome of a game. When a team of $k$ players participates in a game and receives a payout, the objective of Shapley is to distribute the payout among team members in a fair way, i.e., in accordance with their respective contribution. Lundberg and Lee (2017) first proposed the use of Shapley values to explain ML models: each feature may be regarded as a player, the model itself is the game, the prediction of the model is the payout, and the Shapley value for each feature measures its contribution towards the overall prediction. The time complexity of computing Shapley values according to the original formulation is exponential, since it involves examining all possible subsets of features. Lundberg and Lee (2017) formulate approximate versions, viz., Kernel SHAP, Tree SHAP (tree based approximation of SHAP), and Deep SHAP, that may be computed in reasonable time.
5. Evaluating explanations
The subsequent sections describe diverse approaches that have been proposed to explain neural models for text representation/indexing and ranking. In order to measure advances in this area, however, evaluation is possibly the biggest issue that needs to be resolved. Humans constitute the primary target of explanations; the diversity of this target group makes it challenging to formulate an evaluation framework. The consumers of explanations range from lay users, through persons responsible for checking compliance with ethical/legal requirements and domain experts who may not be familiar with the principles underlying IR/NLP models, to IR/NLP researchers and practitioners, whose primary objective may be failure analysis, or to gain insights that guide further development. Of course, this diversity is also true of the users of IR/NLP systems, and the original Cranfield paradigm for evaluating IR had several drawbacks. Some of these have been addressed, while some remain, but over the years, a generally accepted protocol for evaluating IR/NLP systems has emerged. Importantly, a number of large test collections have been created for many different tasks, and their role in advancing the field cannot be overemphasised.
Even though similar, widely accepted protocols and test collections for ExIR are yet to emerge, two broad evaluation paradigms may be identified. The first involves anecdotal evaluation by humans. This seems unavoidable, since most reearch groups do not have the resources to conduct carefully designed, large-scale user studies. While this sort of evaluation may be enough to establish the plausibility of explanation methods, it is not rigorous. The second approach to evaluation utilises different quantitative notions of similarity. When explanations are in the form of simple surrogates that mimic complex models (see §3.4), evaluation is typically based on measuring the similarity between the behaviours of the explanation and the complex model, using standard metrics like rank correlation. Likewise, feature attribution is generally evaluated by measuring the difference in the performance of a model when it is given the original inputs, and when it is given a simplified representation consisting of just the putatively important features. Rationale-based approaches may be evaluated similarly; in cases where human annotated rationales are available, explanations may also be evaluated by measuring the overlap between the generated rationales and the ground truth.
33todo: 3complete this section
6. Document Ranking
Neural models for document ranking are generally categorised as representation-based or interaction-based. Like traditional retrieval models, representation-based models also encode queries and documents as vectors; however, in contrast to the sparse vectors used by traditional approaches, these vectors are dense. On the other hand, interaction-based models capture complex interactions across query-term–document-term pairs. This section provides an overview of efforts to explain both categories of neural ranking methods, but first, we provide a brief review of some foundational attempts to understand traditional IR models (§ 6.1). These early studies were all targeted towards global explanations. Next, in Section 6.2, we discuss the Axiomatic Framework (Fang et al., 2004) — a cornerstone of a substantial body of work on explaining IR — and its recent applications. In Sections 6.36.5, we review other approaches to providing global explanations of Neural Ranking Models (NRMs). Finally, in Sections 6.66.8, we turn to studies devoted to local explanations. Tables 3 and 4 summarise the articles covered in these sections, according to whether they address global or local approaches. We are also preparing a Python package, ir_explain (Saha et al., 2024), that provides implementations of well-known post-hoc explanation strategies (including pointwise, pairwise, and listwise methods) within a single framework.
6.1. Understanding traditional models: the vector space model, and language modeling
Traditional IR systems typically represent documents using terms, i.e., single words (or word stems), and sometimes, additionally, groups of words that have special significance (these are often loosely referred to as phrases), along with their associated weights. The term weights (or feature weights) are calculated using either intuitive heuristics, or on the basis of various mathematical models. Traditional IR models may thus be regarded as explainable by design.
Nonetheless, some early work focused on developing a better understanding of these models. Singhal (1997) carefully examined traditional term-weighting formulae used in the Vector Space Model, explained why they do not work as well as more effective statistical models like BM25, proposed ways to modify the formulae in order to address these flaws, and demonstrated that the modified formulae do indeed yield significant improvements in retrieval effectiveness. Besides conventional text collections, Singhal also considered noisy text collections, such as those generated by running Optical Character Recognition (OCR) on scanned document images, and explained why it is inadvisable to use cosine normalisation for such collections. Kamps et al. (2005) studied similar issues within the context of retrieving XML elements from a collection of XML documents, to explain the importance of length normalisation in XML retrieval. On a related note, Smucker and Allan (2006) compared two well-known smoothing techniques used in Language Modeling, and presented a different explanation for why Dirichlet smoothing outperforms Jelinek-Mercer smoothing, compared to the perspective provided by Zhai and Lafferty (2004). In terms of the criteria introduced in § 3, the explanations provided in (Singhal, 1997; Kamps et al., 2005; Smucker and Allan, 2006) are all model-aware, global explanations. We conclude this section by mentioning I-REX (Roy et al., 2019), an interactive tool that is intended to facilitate further studies along these lines by providing a white-box view into traditional IR models. For example, given a retrieval model, and a query $Q$, the compare command in I-REX provides a pairwise explanation (§ 3.3.1) of the ranks of documents $D_{i},D_{j}$, by displaying the query terms present in $D_{i},D_{j}$ along with their term frequencies, collection frequencies, and overall contribution to the query-document scores.
6.2. The axiomatic framework: improving and explaining IR models
There are certain common principles underlying the different term-weighting formulae that are commonly used, even though they are derived using different mathematical models. In a series of studies (Fang et al., 2004; Fang and Zhai, 2006; Fang et al., 2011), Fang et al. consolidated these principles, and formalised them within an axiomatic framework, as a set of constraints, or axioms, that term-weighting schemes must satisfy, in order to be effective. This framework provides a natural basis for pairwise explanations (§3.3.1), i.e., for reasoning about why a particular document $D_{i}$ is ranked above $D_{j}$ by a ranking model for a given query $Q$. It has been used to explain why certain retrieval models work better than other models, as well as to formulate improved versions of existing retrieval models by better aligning them with IR axioms. (Clinchant and Gaussier, 2009b, a). Later efforts have focused on refining the framework by incorporating additional axioms (Tao and Zhai, 2007; Fang and Zhai, 2006), as well as expanding it to cover pseudo-relevance feedback methods (Clinchant and Gaussier, 2013). Table 2 lists some of the most popular axioms that have been mentioned in the literature. The page at https://www.eecis.udel.edu/~hfang/AX.html is an excellent resource on this topic.
Axiom Group Axiom Name Axiom Statement
TermFrequency TFC1 favour a document with more occurrences of a query term
TFC2 ensure that amount of increase in retrieval score of a document must decrease as more query terms are added
TFC3 favour a document matching more distinct query terms
TDC favour a document matching more discriminating query terms
TF-LNC regulates interaction between term frequency and document length
DocumentLength LNC1 penalize a long document (assuming equal term frequency)
LNC2 over-penalize a long document
SemanticMatches STMC1 (Fang and Zhai, 2006) favour a document containing terms that are more semantically related to query terms
STMC2 (Fang and Zhai, 2006) avoid over-favouring semantically similar terms
STMC3 (Fang and Zhai, 2006) favour a document containing more terms that are semantically related to different query terms
TSFC (Pang et al., 2017) a document having higher exact match should get higher score if their semantic similarity score matches
Term Proximity TP (Tao and Zhai, 2007) favour a document in which query terms appear in closer proximity
Table 2. Popular axioms proposed in the past literature on axiomatic IR research (§ 6.2), adapted from (Fang et al., 2011; Câmara and Hauff, 2020).
The axiomatic framework has been applied in recent times to analyse and understand neural IR models in terms of the common principles underlying traditional IR models (which are, in a sense, explainable by design). Some of these efforts investigate the extent to which various IR models (neural, as well as traditional) satisfy axiomatic constraints. Others try to approximate the ranking generated by a retrieval model via an explainable surrogate, a la LIME (§4.2).
Pang et al. (2017) investigated the difference between the features used in traditional models (BM25 and LMJM) and two pre-BERT NRMs: ARC-I (Hu et al., 2014), a representation based model, and MatchPyramid (Pang et al., 2016b), an interaction based model. Their experiments on the Robust04 and LETOR 4.0 (Qin et al., 2010) datasets suggest that features automatically learned by NRMs are more robust compared to the word-based features used by BM25 and LMJM, in the sense that the performance of NRMs is less affected than that of traditional models when some features are removed. They also confirmed that interaction-based models appear to focus on query related words, while representation-based models seem more concerned with general topic related words, such as those generated by a state-of-the-art LDA based model. Interestingly, Pang et al. noted that ARC-I and MatchPyramid performed poorly on LETOR 4.0 and even worse on Robust04, compared to BM25 and LMJM. Their failure analysis suggests that these NRMs do not satisfy the TDC axiom (Table 2). In addition, due to limitations on memory and time, documents were sometimes truncated to some fixed length, e.g., 500 words. This truncation may reduce the number of query terms present in a document, resulting in a violation of the TFC1 axiom. It was also suggested that semantic matching may be introducing too much noise into the scoring function. A new axiomatic constraint TSFC (Table 2) was proposed, which states that if two documents are of equal length, and have equivalent semantic term matching signals, the document having a larger exact match should be ranked higher. Furthermore, the authors argued that the size of Robust04 is not suitable for training NRMs. 44todo: 4how the words are generated in arc-i pooling e? second version e.
Following the trend of creating diagnostic datasets for NLP and Computer Vision applications, Rennings et al. used selected axioms to construct a similar dataset for IR. Query-document-pair triplets that satisfy TFC1, TFC2, TDC (cf. Table 2) were sampled from the WikiPassageQA (Cohen et al., 2018b) corpus. Artificial documents were constructed to create triplets satisfying LNC2. This diagnostic dataset was used to evaluate the axiomatic efficacy (the fraction of axiomatic constraints that is satisfied by a retrieval model) of both traditional models (BM25 and LMDIR), and neural models like DRMM (Guo et al., 2016), aNMM (Yang et al., 2016), Duet (Mitra et al., 2017) and MatchPyramid (Pang et al., 2016b). The authors observed that statistical models outperform neural models in terms of axiomatic efficacy, but were significantly outperformed by DRMM and aNMM in terms of retrieval effectiveness. In later work, Câmara and Hauff (2020) extended this idea by considering five additional axioms (LNC1, STMC1, STMC2, STMC3, TP). Once again, empirical results showed that DistilBERT (Sanh et al., 2019) significantly outperforms traditional query likelihood methods, but it does not perform as well in terms of axiomatic efficacy. It is particularly counter-intuitive that this observation applies to semantic matching axioms, e.g., STMC1, as well. These results suggest that existing axioms, including the semantic matching axioms, are not adequate for analysing NRMs based on BERT and its variants, and need to be formulated more carefully.
Völske et al. (2021) provided post-hoc explanations of IR systems by constructing an explanation model to approximate the ranking generated by a retrieval model. Recall that each axiom induces relative orderings for certain document pairs; e.g., axiom $A_{i}$ may suggest that $D_{j}$ should be ranked above $D_{k}$. The explanation model (a random forest) may be regarded as a weighted aggregation of these orderings. In the words of the authors, “(a) the explanation model’s parameters reveal the degree to which different axiomatic constraints are important to the retrieval model under consideration, and (b) the fidelity with which the initial ranking can be reconstructed can point to blind spots in the axiom set, which can help to uncover new ranking properties yet to be formalized.” Twenty axioms were used in this study. The behaviour of three statistical models (BM25, TF-IDF, PL2) and five neural models (MP-COS (Pang et al., 2016b), DRMM (Guo et al., 2016), PACRR-DRMM (McDonald et al., 2018), BERT-3S (Akkalyoncu Yilmaz et al., 2019), DAI-MAXP (Dai and Callan, 2019)) were studied using the MSMARCO (Nguyen et al., 2016) and Robust04 collections. [^4] Table 9 summarises information about some datasets commonly used in IR explainability studies. The authors observed that it is difficult to explain the relative ordering of two documents if they have similar retrieval scores, but a big difference in scores can be reliably explained in terms of the axioms. They also noted that explanation effectiveness is comparable for traditional and neural models (in this study, explanation efficacy is measured in terms of how accurately the explanation model can replicate a retrieval model’s ranking), but it was higher for the Robust04 collection, as compared to MSMARCO.
The use of the axiomatic framework in these studies (Rennings et al., ; Câmara and Hauff, 2020; Völske et al., 2021) draws on earlier work by Hagen et al. (2016) that is not directly related to explainability. Hagen et al. were likely the first to formulate relaxed versions of axioms that permitted them to be used for practical analysis,. Axiom TFC1 (cf. Table 2), for example, induces an ordering on a document pair based on the term-frequencies of a query term in the two documents, provided the two documents are of equal length. To be applicable in practice, this equality condition was replaced by an approximate equality constraint: axiom TFC1 was applied to a document pair if the relative difference in the lengths of the two documents was at most 10%. The work discussed above (Rennings et al., ; Câmara and Hauff, 2020; Völske et al., 2021) uses such relaxed formulations of various axioms.
6.3. Global explanations: additive models as explainable surrogates or as inherently explainable models
Sen et al. (2020) studied simple, additive models as surrogates of both statistical and neural retrieval models, by projecting these models onto a 3-dimensional space, with well-understood dimensions, corresponding to term frequency (tf), document frequency (df) and document length (len). For each of the “complex” models, a simple, linear regression model was trained to predict the score of a term in a document using tf, idf and len. For neural models, semantic overlap was introduced as a fourth dimension to capture the contribution of soft (non-exact) term matches to the overall similarity between the query and a document. The regression models are intended to provide linear approximations (as shown in the example below) of the non-linear expressions used by ranking models:
$\mathit{wt}_{\mathcal{M}}(d_{i},D)\approx\alpha\times\mathit{tf}(d_{i},D)~+~\beta\times\mathit{idf}(d_{i})~+~\gamma\times\mathit{len}(D)~+~\delta\times\mathit{semantic\_overlap}(d_{i},D).$
Here, $wt_{\mathcal{M}}(d_{i},D)$ denotes the weight of term $d_{i}$ in a document $D$ as computed by model $\mathcal{M}$, and $\delta=0$ for traditional models. The parameters learned from the regression model were used to “explain” how different ranks were produced by four retrieval models: BM25, LMJM, LMDIR and DRMM.
Generalised Additive Models (GAMs (Hastie and Tibshirani, 1986)) are, as the name suggests, a generalisation of linear regression models. Instead of learning just coefficients, a GAM learns a smooth function for each input feature. The outputs of these functions are simply added together to obtain the prediction $y$ for an input instance. GAMs are regarded as interpretable (Lou et al., 2012). Zhuang et al. (2021) applied GAMs to the Learning to Ranking (LETOR) problem. In their algorithm, NeuralRankGAM, each function constituting the GAM was represented by a simple feed forward neural network, leading to the claim that the resultant architecture is inherently interpretable. The authors adapted piece-wise regression (Muggeo, 2003) to construct simpler, distilled versions of these networks in order to make inferencing faster, as well as to make the models potentially more interpretable. Of course, it is debatable whether a “simple” model involving “only” a few thousand parameters is actually interpretable. On three LETOR datasets (Yahoo (Chapelle and Chang, 2011), Web30K (Qin and Liu, 2013), and a private dataset sub-sampled from the Chrome Web Store logs), NeuralRankGAM outperformed, in terms of NDCG, other GAMs (TreeGAM (Lou et al., 2012) and Tree RankGAM (Ke et al., 2017)) that were used as baselines. Interpretability was studied using visualisations of occlusion-based feature attribution scores (§ 4.3). The model appeared no more interpretable than its competitors.
Lucchese et al. (2022) proposed another “inherently interpretable” ranking model based on LambdaMART (Burges et al., 2011), that has a similar structure as GAMs, but uses ensembles of trees. Their approach uses a modified boosting LambdaMART algorithm in three stages. In the first stage, an ensemble of trees is used, with each tree working on a single feature. The interaction between distinct pairs of features (as identified in the first stage) is modeled in the second stage. In the last stage, the LambdaMART algorithm is restricted to use only the top $k$ feature pairs obtained from the previous stages. The model was compared against NeuralRankGAM (discussed above) and a pointwise ranking algorithm EBM (Lou et al., 2013), on three learning to rank datasets, Istella-S (Lucchese et al., 2016), Web30K, and Yahoo. The proposed model resulted in further improvements in NDCG, but its claim of intepretability is open to debate.
6.4. Global explanations based on probing
Recall that probing (Section 4.1) attempts to characterise linguistic features that are encoded within the representations generated at intermediate layers of a pre-trained network. In an IR context, early experiments with probing were reported by Palangi et al. (2016) and Cohen et al. (2018a). Cohen et al. trained a multilayered LSTM network for an IR-oriented task, viz., answer passage retrieval for the CQA (Surdeanu et al., 2008) and WikiQA (Yang et al., 2015) datasets. After the weights were frozen, the probing framework proposed in (Alain and Bengio, 2017) was applied to each layer of the network. The representations were used for NLP tasks such as POS tagging on the Penn TreeBank (Marcus et al., 1994), NER on CoNLL-2003 (Tjong Kim Sang and De Meulder, 2003), sentiment classification on the IMDB review collection (Maas et al., 2011), and textual entailment for Stanford’s Natural Language Inference (SNLI) corpus (Bowman et al., 2015). Cohen et al. found that initial layers appear to capture NER and POS features, while the later (second LSTM) layer is concerned with higher-level, abstract topical representations. Interestingly, the study by Palangi et al. indicates that such conclusions may be collection-dependent, i.e., the abstract representations appear to capture different features for different collections. For example, representations from the WikiQA model are more effective for POS tagging and NER than those from the CQA model, but they perform poorly on the sentiment classification and entailment tasks.
Qiao et al. (2019) highlighted the interaction-based, sequence-to-sequence nature of BERT using a few different BERT-based rankers. A representation-based BERT model was designed by taking the individual [CLS] embeddings for both queries and documents, and computing the cosine similarity between them. Experiments on the MSMARCO passage reranking and TREC Web Track document ranking tasks showed that, on both datasets, the representation-based BERT model performed poorly. This suggests that BERT should not be regarded as a representation-based model. Interestingly, Qiao et al. observed that, while monoBERT (Nogueira and Cho, 2019) outperformed the then-state-of-the-art NRMs (K-NRM, Conv-KNRM) on MSMARCO, it failed to perform well on document ranking. They concluded that passage ranking for MSMARCO is more related to a question answering (QA) task, and therefore, better suited for a seq2seq model. Indeed, MSMARCO is basically a QA dataset, where precise matches are important. In contrast, traditional IR datasets, where recall is also important were used to train Conv-KNRM. It would be interesting to fine-tune BERT on such datasets, and then compare its performance against Conv-KNRM. On the other hand, their experiments showed that removing a single term from a document affects its score as calculated by the Conv-KNRM model, but monoBERT is more robust to such changes. Qiao et al. also contributed to the ongoing debate about attention (this is summarised in Section 9.1). They observed that the [CLS] and [SEP] tokens received the most attention. Further, even though removing stopwords does not affect the effectiveness of a ranker, stopwords and non-stop words received similar attention scores.
A more recent series of experiments by MacAvaney et al. (2022) uses a different form of probing to analyse the sensitivity of NRMs to various changes in textual properties. In a sense, MacAvaney et al.’s approach is similar in spirit to feature-attribution studies: it looks at how the output changes when changes are made to input features. However, their objective is to obtain a global characterisation of NRMs, rather than determine which features of a particular input instance are important. A given NRM (trained on the MSMARCO passage ranking task) was first probed using a query-document pair $\left\langle Q,D\right\rangle$, and then using $\left\langle Q,D^{\prime}\right\rangle$, with $D^{\prime}$ being a perturbation of $D$. Four broad classes of perturbations were studied:  (a) “Measure and Match”, in which $D^{\prime}$ was generated by varying a specific attribute of $D$ (e.g. term frequency), while keeping other attributes (e.g., document length) fixed;  (b) simple textual manipulation (changing the order of words in a passage, introducing typographic errors, etc.);  (c) “Dataset Transfer Probes” (DTPs), which involve more complex, stylistic changes, e.g., changes in the level of fluency, formality, or the succinctness of a passage; and  (d) changes in factual content, which typically involve replacing one entity (e.g., New York) by another entity of the same type (e.g., Moscow). For the first two types of perturbations, synthetic documents were created, while DTPs were created from datasets specifically designed for the purpose (e.g., summarisation datasets). If a perturbation resulted in significantly different scores being computed for $D$ and $D^{\prime}$, the corresponding NRM was regarded as being sensitive to the perturbation. Experiments were conducted on the TREC-DL 19 passage ranking data and ANTIQUE (Hashemi et al., 2020). A wide range of models were studied: BM25, WMD (Kusner et al., 2015), SBERT (Reimers and Gurevych, 2019), LGBM (Ke et al., 2017), VBERT (Nogueira and Cho, 2019), T5 (Raffel et al., 2020), EPIC (MacAvaney et al., 2020), DT5Q (Nogueira and Lin, 2019), ANCE (Xiong et al., 2021b), and ColBERT (Khattab and Zaharia, 2020). Empirical observations include the following.
  • Traditional models are better than NRMs at handling typographical errors artificially introduced into the text.
  • When non-relevant content is added to a document, its BM25 score generally decreases, but the score computed by contextualized NRMs often increases.
  • Most contextualized models work well on full texts, while BM25 has a strong affinity towards summaries.
  • Some neural models, viz., ColBERT, T5 and WMD, appear to “memorise” facts from the corpora they are trained on, and generally assign lower scores to factually incorrect passages (e.g., “New York is the capital of Russia”).
6.5. Global explanations for BERT-based NRMs: what knowledge do they possess?
Instead of probing, some studies have examined attention maps and query-document interactions, across different layers and heads, to understand what sort of knowledge BERT-based models encode internally that other statistical and older NRMs do not.
Zhan et al. (2020) studied the monoBERT model focusing on both its attention mechanisms and query-document interactions. Once again, the MSMARCO passage ranking task was used for experiments. As in (Qiao et al., 2019) (discussed above) and similar studies conducted within the NLP community (Clark et al., 2019; Rogers et al., 2020; Kovaleva et al., 2019), empirical results show that BERT puts redundant attention weights on [CLS] and [SEP] tokens. Query-document interactions were found to play a major role in document ranking. Their observations regarding the behaviour of different layers of BERT echo earlier studies: interactions are not important in the initial layers of BERT, where it appears to learns more context specific information; in the later layers, it heavily uses interactions between a query and document to predict relevance. These findings were confirmed by experiments with probing (§ 4.1) and attribution (§ 4.3) techniques.
Recently, ColBERT has gained popularity in the IR community. Its scoring function is as follows:
$\mathit{Score}_{\mathrm{ColBERT}}(Q,D)=\sum_{q_{i}\in Q}\max_{d_{j}\in D}E_{q_{i}}E_{d_{j}}^{T}$
where $E_{q_{i}}$ and $E_{d_{j}}$ denote the embeddings of the $i$-th query term and $j$-th document token, respectively. This scoring function is superficially similar to early embedding-based similarity functions, except that $E_{q_{i}}$ and $E_{d_{j}}$ are heavily contextualised representations of query / document tokens. In earlier neural IR models, like DRMM (Guo et al., 2016), term heuristics (e.g., inverse document frequency) were explicitly involved in the architecture; ColBERT eliminates the need for such measures. Formal et al. (2021) present a “white-box analysis” of ColBERT on the TREC-DL 2019 and 2020 passage retrieval tasks. For each query term $q_{i}$, the authors analysed the distributions of exact matches and soft (non-exact) matches between $q_{i}$ and document tokens. They observed that exact matches are more important for high-IDF query terms. This is not unexpected, since such terms are likely to be specific, and therefore unlikely to have high semantic overlap with other terms in the document. The authors also estimated the impact of each $q_{i}$ on ColBERT’s ranking using a leave-one-out study: if the correlation ($\tau_{\mathit{AP}}$ (Yilmaz et al., 2008)) between the ranked lists for $Q$ and $Q\setminus\{q_{i}\}$ is high (close to 1, say), then $q_{i}$ is not important. Experiments show that ColBERT gives more importance to higher IDF terms, as expected.
Approach Explanation type Dataset(s) Codebase
Fang et al. (2004) Proposed axiomatic framework TREC678 NA
Hagen et al. (2016) Relaxed axioms and used for re-ranking Clueweb09b github.com/webis-de/ir_axioms
Pang et al. (2017) Finding interaction vs representation focused NRMs Robust, LETOR 4.0 NA
Rennings et al. Diagnostic datasets with axioms Diagnostic dataset created from WikiPassageQA NA
Câmara and Hauff (2020) Diagnostic datasets with axioms Diagnostic dataset (proposed) github.com/ArthurCamara/bert-axioms
Sen et al. (2020) Additive models MSMARCO passage NA
Zhuang et al. (2021) GAMs EB30K, YAHOO, Chrome Web Store (private data) github.com/tensorflow/ranking/releases/tag/v0.3.1
Lucchese et al. (2022) Ensemble of trees WEB30K, YAHOO, Istella-S github.com/veneres/ilmart
Cohen et al. (2018a) Probing complex rankers CQA, WikiQA NA
Qiao et al. (2019) Understanding BERT for ranking MSMARCO passage, Clueweb09b NA
MacAvaney et al. (2022) Probing NRMs MSMARCO passage and created MMPs, TMPs from TREC-DL 2019 and MSMARCO passage github.com/allenai/abnirml
Formal et al. (2021) Understanding ColBERT TREC-DL 2019, 2020 passage NA
Table 3. List of global ExIR methods covered in Sections 6.26.5, along with datasets used, and links to codebases, where available.
6.6. Local explanations: LIME and SHAP inspired approaches
We now turn to studies that propose methods for constructing local explanations. As discussed in Section 4.2, LIME based local, post-hoc explanations have become popular for classification tasks. Following (Ribeiro et al., 2016), Verma and Ganguly (2019) proposed LIRME, an adaptation of LIME that provides pointwise explanations for document ranking. $\mathcal{M}_{C}(Q,D)$, the score that an IR model $\mathcal{M}_{C}$ computes for a query-document pair $(Q,D)$, is explained by a vector of weights corresponding to different tokens, where the weight for a token indicates its contribution to $\mathcal{M}_{C}(Q,D)$. To estimate this explanation vector, perturbed versions of $D$ were generated by selecting subsets of words from $D$, and the effects of the perturbations on the retrieval score assigned by $\mathcal{M}_{C}$ were noted. Experiments were conducted on the TREC-8 adhoc collection. The authors also proposed two metrics to evaluate the generated explanations: consistency and correctness. Consistency measures how stable the explanation vector is across perturbations, while correctness quantifies whether an explanation vector assigns higher weights to “relevant” terms. Perhaps the biggest flaw of LIRME is the overly simplistic definition of ground truth about which terms are actually relevant: this is determined based on the weights of terms occurring in relevant documents, calculated using a simple statistical model (LM-JM).
Singh and Anand (2019) also incorporated ideas from LIME to create pointwise explanations. Their method, EXS, uses a linear SVM as the simple model that approximates an arbitrary, black-box ranker. As usual, perturbed versions of a document $D$ are constructed by deleting random words from random positions in $D$. The authors mainly focus on three different ways to estimate the relevance of a perturbed instance of $D$:  (a) top-k binary: the perturbed version is regarded as relevant if its score, as computed by the simple model (linear SVM), is more than that of the $k^{th}$ document in the original ranked list, and is regarded as non-relevant otherwise;  (b) rank based: this is similar to the above strategy, except that the relevance score for a relevant document decays with its rank, instead of being 1; and  (c) score based: In this method, relevance scores are real-valued, rather than binary and are determined by the relative difference in score between the perturbed document and the top-ranked document. The form of an explanation is similar to that for LIRME: it consists of a graph displaying which words are strong indicators of relevance for a particular query. EXS was used to explain neural retrieval models such as DRMM (Guo et al., 2016) and DESM (Nalisnick et al., 2016), on traditional news collections used at TREC (AP, LATIMES, Robust04 and Financial Times).
Fernando et al. (2019) explored the use of Shapley values (§ 4.3.2) for model-introspective explanations for IR. They used DeepSHAP, which internally computes the Shapley values of terms present in a document (a high value indicates a bigger contribution to the prediction of relevance). The authors compared various approaches for creating a reference document, in order to determine which reference document provides the best explanations for retrieval models. For example, reference documents were constructed by averaging the embeddings of either all words in the vocabulary or a set of words randomly sampled from the collection, excluding the top 1000 documents retrieved by the query. These explanation strategies were investigated on three NRMs: DRMM, MatchPyramid (Pang et al., 2016b), and PACRR-DRMM (McDonald et al., 2018). LIME based explanations were used as a baseline. Although these models use standard neural network architectures, empirical results showed that no single reference document yields the best explanations across different neural ranking models. Further, the performance of these retrieval models is also sensitive with respect to different reference inputs.
6.7. Extracting and evaluating alignment rationales as local explanations
As discused in Section 3.4.2, a rationale is a set of tokens in an input $I$ that best explains a model’s decision for $I$. In an IR context, alignment rationales, which establish a correspondence between portions (or spans) of a query and a document, are particularly useful. Consider the hypothetical query Where is SIGIR 2022? (Kim et al., 2022), in response to which a neural model may plausibly retrieve the span “SIGIR 2022 will be held in Madrid”. The underlined parts of the query and the document constitute an alignment rationale (AR), since Madrid, a location, is a response to the Where is portion of the query. [^5] Clearly, for conventional IR models, alignment rationales are limited to term matches between the query and the document. Kim et al. (2022) proposed various perturbation-based metrics to evaluate how faithfully ARs reflect a model’s behaviour. Deletions and substitutions were considered as possible perturbations. Empirical findings with a BERT-based ranker (DAI-MAX (Dai and Callan, 2019)) that was trained on the MSMARCO document ranking task suggest that metrics based on substitutions are more reasonable compared to those based on deletions for measuring faithfulness.
Leonhardt et al. (2023) proposed a rationale-based ‘interpretable’ ranking algorithm. The top 20 sentences were extracted from a document as a rationale; a ‘sparse’ representation consisting only of these sentences was used to rank each document. Two sentence-scoring approaches were explored:  (a) an end-to-end gradient-based optimisation technique that chooses sentences in order to maximise retrieval effectiveness in the ranking phase; and  (b) choosing sentences deterministically, e.g., based on their tf-idf scores with respect to the query. For ranking, a monoBERT-style architecture was used with the query and the document rationales. The system’s retrieval effectiveness on the TREC-DL, Core 17, Clueweb09, and BEIR (Thakur et al., 2021) datasets was at par with state-of-the-art models (BERT-3S (Akkalyoncu Yilmaz et al., 2019), Doc-labeled (Dai and Callan, 2019) (DAI-MAXP), BERT-CLS (Nogueira and Cho, 2019) (monoBERT)). Additionally, the quality of the rationales was evaluated by humans.
Evaluating rationales
Pandian et al. (2024) proposed that the quality of rationales be evaluated using intrinsic and extrinsic measures. Intrinsic evaluation quantifies the ‘consistency’ (correlation) between the ranks of the top-$k$ documents when scores are computed using the full-text vs. only rationales (specifically, the concatenation of the top $m$ rationales, where rationale scores were computed using occlusion). [^6] In occlusion (§4.3), document segments are masked in turn; the resultant changes in the document’s score indicate the importance of the respective segment. On the other hand, extrinsic evaluation quantifies how well the top $m$ rationales overlap with the relevant passages of the documents. Relevant passages were identified by aligning the MSMARCO passage and document datasets via the “QnA” version of MSMARCO. Experiments with these datasets suggest that the retrieval effectiveness of a ranking model (both BM25 and NRMs e.g., ColBERT) does not correlate with the proposed intrinsic or extrinsic effectiveness measures.
6.8. Using query expansion for local, listwise explanations
Query expansion (QE) based approaches (see § 3.4.1) approximate a complex retrieval algorithm with a simpler one, by adding ‘explanation’ terms to the original query. IntentEXS (Singh and Anand, 2020) and Multiplex (Lyu and Anand, 2023) are both examples of such approaches, and broadly follow a common framework. A ranked list is viewed as a set of preference pairs of documents. The goal is to identify a set of explanation terms to add to the original query, such that a simple IR model (e.g., BM25) ranks most preference pairs in the same order as a given NRM. While IntentEXS uses a greedy preference pair coverage algorithm to maximise the above objective, Multiplex uses the modern Geno method (Laue et al., 2019) for optimisation.
IntentEXS used LMJM as the simple ranker to explain three pre-BERT neural models: DRMM (Guo et al., 2016), P-DRMM (McDonald et al., 2018), M-Pyramid (Pang et al., 2016a). The effectiveness of the explanation terms was evaluated using two metrics. Fidelity, or predictive accuracy, was measured using the correlation (Kendall’s $\tau$) between the ranked lists of documents retrieved by LMJM and the NRMs. Next, expansion terms generated using RM3 (Lavrenko and Croft, 2001), EMB (Diaz et al., 2016) or DESM (Nalisnick et al., 2016) were regarded as ground truth, and descriptive accuracy was measured in terms of the overlap between the explanation terms and this ‘ground truth’. Experiments were conducted on Clueweb09b (Callan et al., 2009). When the entire ranked list was considered, Kendall’s $\tau$ was found to be <0.5<0.5 for the three NRMs. When only the top $10$ documents were considered, a $\tau$ greater than $0.5$ $(0.537)$ was observed for only P-DRMM. Given these low $\tau$ values, this method does not seem to adequately explain the workings of NRMs. The authors showed, however, that the framework can be used to identify model and temporal biases. For instance, in response to the query ‘fidel castro and electoral college 2008 results’, DRMM considers documents related to events in 2004 or later. Similarly, for the query ‘how to find mean’, DRMM correctly captures the statistical intent behind the query, whereas DESM fails to do so. Multiplex, on the other hand, attempted to explain three NRMs (DRMM, monoBERT and DPR (Karpukhin et al., 2020)) using a combination of three simple rankers: i) term matching (BM25), ii) position aware, and iii) semantic similarity. A fidelity measure (the extent to which preference pairs are in agreement with the original ranking) was the evaluation metric. Not all pairs were considered, however; evaluation was limited to a random sample of preference pairs for which the rank difference was large. Within this restricted setup, Multiplex achieved a high fidelity score on Clueweb09b and TREC-DL 2019.
Approach Explanation type Dataset(s) Codebase
LIRME (Verma and Ganguly, 2019) Adaptations of LIME TREC-8 github.com/gdebasis/irexplain
EXS (Singh and Anand, 2019) TREC678, Robust github.com/GarfieldLyu/EXS
Fernando et al. (2019) SHAP-based Robust No URL provided
Kim et al. (2022) Evaluating rationale faithfulness MSMARCO docs; TREC-DL 2019 queries github.com/youngwoo-umass/alignment_rationale
Leonhardt et al. (2023) Ranking based on rationales TREC-DL 2019, Clueweb09b, Core 17, BEIR benchmark No URL provided
Pandian et al. (2024) Extracting + evaluating rationales MSMARCO (pass. & docs); TREC-DL 2020 queries github.com/saranpandian/XAIR-evaluation-metric
Singh and Anand (2020) Explanation terms / expanded queries Clueweb09b github.com/GarfieldLyu/EXS
Multiplex (Lyu and Anand, 2023) Clueweb09b, MSMARCO passage and TREC-DL 2019 queries github.com/GarfieldLyu/RankingExplanation
Yu et al. (2022) Diversity-aware listwise explanation Wiki, MIMICS github.com/PxYu/LiEGe-SIGIR2022
BFS (Llordes et al., 2023) Listwise explanation MSMARCO passage and TREC-DL 2019 queries github.com/micllordes/eliBM25
Table 4. List of local ExIR methods covered in Sections 6.66.8, along with datasets used, and links to codebases, where available.
Yu et al. (2022) also consider the problem of providing a listwise explanation of documents. Strictly speaking, however, their objective is not to explain why a given list of documents was retrieved, or ranked in the given order; it would be more appropriate to term their “explanations” as annotations. These annotations are intended to make it easier for a user to consume the information contained in the list of documents. Yu et al. provide two types of annotations. Given a (relatively broad) query involving multiple aspects, the first type of annotation “explains” which aspects of the query are covered by each document in the ranked list. The second type of annotation highlights the novel aspects covered by a particular document, i.e., the aspects that are not already covered by documents ranked above it. The annotations were created by modifying the encoder and decoder components of a Transformer network. Specifically, inter-document and intra-document attention layers were added.
Experiments were conducted on a newly constructed Wikipedia corpus, and the MIMICS (Zamani et al., 2020) dataset. The title of a Wikipedia article was considered as a query. Each section of the article was treated as a unit of retrieval (i.e., as a single document). The heading of any section was treated as the ground-truth / gold-standard explanation that specifies which aspects of the query are covered by the corresponding document. However, as Wikipedia articles are topically organised, most sections cover only a single aspect of a query. Therefore, documents covering multiple, overlapping query aspects were artificially created. Given a Wikipedia article, three sections were randomly sampled from it (without replacement) at a time. Two documents were created from these three sections such that the two documents had one section in common. Thus, from an article containing $n$ sections, $\lfloor n/3\rfloor\times 2$ documents were created. The second dataset (MIMICS) is a search clarification dataset with the search queries sampled from Bing query logs. For each query, the clarifying questions are considered as query aspects.
A group of models from search result explanation tasks, TextRank (Mihalcea and Tarau, ), a topic sensitive PageRank variant of TextRank, LIME based explanations for monoBERT, KeyBERT (Grootendorst, 2020), KeyBERT with maximal marginal relevance (Carbonell and Goldstein, 1998), GenEx (Rahimi et al., 2021), HistGen (Zhang et al., 2019), NMIR (Hashemi et al., 2021) and BART (Lewis et al., 2020) were used as baselines. The quality of explanations was evaluated with BLEU and ROUGE scores. Additionally, BERTScore (Zhang et al., 2020) was computed to measure the semantic similarity between the generated explanations and ground truth explanations. A divergence based metric, Div, was also employed. The proposed architecture outperforms the baselines in terms of the above metrics. 55todo: 5cut down kora jay, too much lekha hoyechge
Llordes et al. (2023) proposed a more traditional, QE-based explanation algorithm, BFS, which uses Jaccard and rank-biased overlap (RBO (Webber et al., 2010)) to quantify the similarity between the ranked lists produced by an NRM and another, produced using the expanded query by traditional ranker, such as BM25. BFS is a heuristic search algorithm that explores the best set of terms sampled from pseudo-relevance feedback from documents. Intuitively, if adding a term increases the overlap between the two ranked lists, it explores that combination. Experimental results on TREC-DL 2019 data for popular NRMs, such as ColBERT and ANCE (Xiong et al., 2021a), were modest, with BFS achieving a maximum RBO of around 0.5 for the top 10 documents. Additionally, the running time of BFS is also high.
7. Understanding Embeddings
Document ranking methods optimise relevance by relying on text representations that capture semantic meaning, with word embeddings playing a crucial role by providing dense, low-dimensional vectors encoding linguistic and contextual relationships. Word embeddings gained widespread acceptance after pre-trained embeddings, e.g., Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014), became available. Their use as input to complex neural models is now ubiquitous in NLP/IR. Understanding embeddings is thus key to explaining NLP and IR methods, as they underpin how models interpret and process text. While post-hoc explanations attempt to explain how embeddings capture semantics, a straightforward interpretation of individual embedding dimensions remains an open problem. A detailed review of research on explaining word embeddings can be found in (Zini and Awad, 2022). This section supplements (Zini and Awad, 2022) by covering recent related work. See Table 5 for a summary of the articles discussed here.
7.1. Transformation of embedding spaces
A general approach to interpreting embeddings involves applying a transformation to every word to project it into a space with interpretable dimensions. The survey by Zini and Awad (2022, Section 5.2) describes some of these techniques. Below, we summarise Densifier (Rothe et al., 2016), a variant of which is described in (Zini and Awad, 2022), and discuss a recent study (Mathew et al., 2020) that was not covered by earlier surveys.
Let $E\in\mathbb{R}^{n\times d}$ denote the embedding matrix, i.e., the stack of $n$ embedding vectors of size $d$, where $n$ is the number of words in the vocabulary. The objective of Densifier (Rothe et al., 2016) is to look for an orthogonal matrix $Q$ such that the product $EQ$ is “interpretable” in the sense that the values of its first $k$ dimensions have a good correlation with linguistic features. More precisely, suppose $v$ and $w$ are two words and let $l$ be a linguistic feature. Intuitively, the objective function of Densifier looks for a unit vector $q_{(l)}$ such that $\|q_{(l)}^{T}(v-w)\|$ is high (resp., low) when $v$ and $w$ differ (resp., agree) with regard to $l$. In a later work, Dufter and Schütze (2019) proposed a small modification to the objective function that allows them to obtain a closed-form analytic solution for $Q$ that is hyper-parameter free. The transformed embeddings work well for lexicon induction and word analogy tasks. They used a simple linear SVM and Densifier as interpretability baselines. Both (Rothe et al., 2016) and (Dufter and Schütze, 2019) used the SemEval2015 Task 10B and Czech movie reviews dataset; additionally, Dufter and Schütze (2019) used the Word Analogy dataset. Mathew et al. (2020) proposed POLAR, a framework for ensuring the interpretability of pre-trained embedding vectors. They take a set of $m$ polar opposite antonym pairs. Suppose $(w_{i},w^{\prime}_{i})$ is one such antonym pair, and let $v_{i}$, $v^{\prime}_{i}$ be their corresponding pre-trained embeddings. A new basis $\mathcal{B}$ is constructed from the vectors $v_{i}-v^{\prime}_{i}$ for the $m$ pairs. If $x_{d}$ is the original pre-trained embedding of a word, then a standard transformation can be used to compute its projection to the subspace spanned by $\mathcal{B}$. POLAR (Mathew et al., 2020) was used to transform and evaluate Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) based pre-trained word embeddings on a wide range of datasets: News Classification, Noun Phrase Bracketing, Question Classification, Capturing Discriminative Attributes, Word Analogy, Sentiment Classification, and Word Similarity tasks. Across most of the tasks, POLAR makes the vectors interpretable while retaining performance. Interpretability was measured by comparing the top 5 dimensions of words with Word2Vec and human annotators.
7.2. Changing the objective function of word embeddings
Modifying the embedding objective function is another approach to enhance interpretability. This involves incorporating components that preserve the underlying semantic structure while aligning embedding dimensions with predefined concepts. Senel et al. (2021) modified GloVe’s objective function to embed concepts within vector dimensions, using external concept word resources. Their cost function weighted embedding dimensions based on concept group membership, encouraging alignment with the corresponding semantic space. The embedding dimension size matched the number of concept word groups (300 from Roget’s Thesaurus). However, this approach did not utilize negative vector directions. In follow-up work, Senel et al. (2022) incorporated both positive and negative vector directions by modifying GloVe and Word2Vec loss functions. The modified cost function included components for both positive and negative concepts, assigning higher/negative weights based on positive/negative concept group membership (e.g., “good” and “bad” as opposites). Evaluations included Word Similarity and Analogy tests, along with sentiment analysis, question classification, and news classification, using SVMs trained on average input text embeddings.
7.3. The power of contextual embedding
Pre-trained embeddings, such as Word2Vec and GloVe, offer context-independent word representations, which is a limitation for polysems. For example, “apple” has the same representation in both sentences: “Apple has great health benefits” and “Apple has released a new iPad”. Context-dependent embeddings, obtained by processing entire text through models, such as CoVe (McCann et al., 2017), ELMo (Peters et al., 2018), and BERT (Devlin et al., 2019), address this issue.
Tenney et al. (2019b) investigated the advantage of contextualized embeddings by probing spans of contextual embedding vectors, concatenating them and passing them through a multi-layered network. Using CoVe, ELMo, GPT (Radford and Narasimhan, 2018), and BERT, they probed for sub-sentence structure. They used CoVe’s GloVe vectors and ELMo’s initial character-CNN activations as context-independent baselines. Experiments across various NLP tasks (part-of-speech tagging, constituent labeling, dependency parsing, named entity recognition, semantic role labeling, coreference resolution, semantic proto-role labeling, and relation classification) showed that contextualized embeddings capture syntax and long-range linguistic information better than non-contextualized ones, as measured by micro-averaged F1 scores. A randomized version of ELMo (with weights being replaced by random orthonormal matrices) performed reasonably well, but the full ELMo achieved nearly 70% improvement, indicating ELMo’s architecture plays a significant role, though its precise nature remains unexplained.
Approach Explanation type Codebase
Densifier (Rothe et al., 2016) Transforming word embeddings No URL provided
Dufter and Schütze (2019) github.com/pdufter/densray
POLAR (Mathew et al., 2020) Transforming word embeddings to interpretable antonyms github.com/Sandipan99/POLAR
Senel et al. (2021) Concept (positive direction) based embeddings github.com/koc-lab/imparting-interpretability
Senel et al. (2022) Concept (positive and negative directions) based embeddings No URL provided
Tenney et al. (2019b) Contextual vs non-contextual embeddings over their architecture github.com/nyu-mll/jiant
Table 5. Summary of explanations for embedding vectors (Section 7).
8. Sequence Models
In this section, we discuss interpretability of sequence models, focusing on specific post-hoc explanation approaches for LSTM (Hochreiter and Schmidhuber, 1997) and hierarchical explanations. For a comprehensive discussion, refer to Zini and Awad (2022).
8.1. Predictive power of individual words
Text classification can often be explained through key words, e.g., for intuitive tasks such as sentiment analysis, the word “delighted” usually indicates positive sentiment. Murdoch and Szlam (2017) analyze LSTM outputs by measuring the influence of individual words. Let the activation after processing the $t$-th word be defined as $h_{t}=o_{t}\odot\mathrm{tanh}(c_{t})$, where $c_{t}$ is the memory state, $o_{t}$ is the output gate and $\odot$ is the element-wise multiplication operator. After processing the full sequence of length $T$, the classification probability of the $i$-th class (among $C$ classes) is given by the softmax $p_{i}=\mathrm{exp}(W_{i}h_{T})\big{/}\sum_{j=1}^{C}\mathrm{exp}(W_{j}h_{T})$, where $W$ is an appropriate parameter matrix and $W_{i}$ denotes the $i$-th row of $W$. Murdoch and Szlam (2017) define the contribution of the $j$-th word towards the $i$-th class using the incremental difference of activation vectors passed through the output gate, as $\beta_{i,j}=\mathrm{exp}\left((W_{i}(o_{T}\odot(\mathrm{tanh}(c_{j})-\mathrm{tanh}(c_{j-1})))\right)$. Subsequently, the output of an LSTM network can be decomposed as
(1) $\mathrm{exp}(W_{i}h_{T})=\mathrm{exp}\left(\sum_{j=1}^{T}W_{i}(o_{T}\odot(\mathrm{tanh}(c_{j})-\mathrm{tanh}(c_{j-1})))\right)=\prod_{j=1}^{T}\beta_{i,j}$
In a similar fashion, they define $\gamma_{i,j}$, in terms of the forget gates to reflect the upstream changes made to $c_{j}$ after processing the $j$-th word. The phrases are scored and ranked using above $\beta_{i,j}$ and $\gamma_{i,j}$ values by averaging over all the documents. The predictive score (applying $\beta_{i,j}$ and $\gamma_{i,j}$ individually) $\mathrm{ps}_{i}$ for a class $i$ of a phrase is calculated as the ratio of average contribution of that phrase to the prediction of class $i$ to class $j$ across all occurrences of that phrase.
For experimental evaluation (binary classification), phrases were tagged with the label of the class with the maximum predictive score ($\mathrm{ps}$). A rule-based classifier, approximating LSTM output, performs string-based pattern matching. Given a document and a list of phrases sorted by $\mathrm{ps}$, it searches for each phrase, returning its label upon a match and stopping after the first match.
Experiments used the SST and Yelp datasets (sentiment analysis) and the Wiki Movies dataset (question answering) (Murdoch and Szlam, 2017). For question answering, a question is encoded with an LSTM, augmented with document word embeddings, and fed to another LSTM with a softmax classifier to predict answer entities. Quantitative evaluations show the rule-based classifier can explain LSTM output via key phrases, though approximation errors exist. For example, while the LSTM correctly classified a sentence like “Still, it gets the job done – a sleepy afternoon rental” as negative, the pattern matcher tagged “gets the job done” as positive, suggesting the LSTM uses broader contextual information.
In a later work, Murdoch et al. (2018) proposed contextual decomposition to address limitations of their prior work (Murdoch and Szlam, 2017), which failed to identify phrases like “used to be” as strongly negative within a sentence. The method decomposes LSTM output and cells into phrase-specific ($\beta$) and other-factor ($\gamma$) contributions, such that the activation and memory states are represented as $h_{t}=\beta_{t}+\gamma_{t}$ and $c_{t}=\beta_{t}^{c}+\gamma_{t}^{c}$ respectively. Recursive computation groups terms derived from the phrase and external factors, identifying sentiment interactions. Contextual decomposition identifies negations and captures LSTM prediction dynamics. For example, it correctly identifies “used to be” as strongly negative and “my favorite” as positive in “used to be my favorite,” shown via per-word $\beta$ heatmaps. This approach was compared to cell decomposition (Murdoch and Szlam, 2017), integrated gradients (Sundararajan and Najmi, 2020), leave-one-out (Li et al., 2016), and gradient times input through experiments conducted on Yelp and SST (binary sentiment analysis), by investigating phrase-level and word-wise CD scores.
8.2. Feature interactions in a hierarchical structure
Some follow-up works aimed to identify and hierarchically use groups of the most predictive features (words), focusing on adjacent words. Singh et al. (2019) extended contextual decomposition (CD) to generic deep networks, generalizing Murdoch et al. (2018)’s cell/output decomposition to layer-wise contextual decomposition:
$g^{CD}(x)=g^{CD}_{L}(g^{CD}_{L-1}(\ldots(g^{CD}_{2}(g^{CD}_{1}(x)))))$
where, $x$ is the input to the network and $L$ is the number of layers in the neural network. On each layer $l$, $g^{CD}_{l}(x)$ is composed of two components: $\beta_{l}(x)$, measuring the importance of the feature group present in the input $x$, and $\gamma_{l}(x)$, which captures the contributions of rest of the tokens in the input. For each layer, $g^{CD}_{l}(x)=\beta_{l}(x)+\gamma_{l}(x)$. One can measure the score of $\beta_{l}(x)$ and $\gamma_{l}(x)$ values on each layer (from the discussion in the earlier section (Murdoch et al., 2018)).
The proposed method generalizes to CNNs, with contextual decompositions for convolution, max-pooling, dropout, and ReLU. Agglomerative contextual decomposition (ACD) was proposed to combine word groups, resembling agglomerative clustering. It combines contextual decomposition scores of each feature, merging adjacent features bottom-up via a priority queue containing words, phrases (or pixels), and their contribution scores. ACD, using SST, MNIST, and ImageNet, identifies incorrect predictions from phrase interactions. ACD is robust to adversarial perturbations (small noise added to inputs), producing similar hierarchical explanations for perturbed and original instances, capturing essential input parts.
(a) HEDGE method used by Chen et al. (2020).
(b) ACD method used by Singh et al. (2019).
Figure 1. Reproducing from  (Chen et al., 2020), showing the difference in interpretability results between HEDGE and ACD (Singh et al., 2019). LSTM makes a wrong prediction as it missed the interaction between never and fails.
HEDGE (Chen et al., 2020) offers another hierarchical explanation. Investigating LSTM, CNN, and BERT on SST and IMDB, HEDGE iteratively divides text chunks into two segments, involving two optimizations: segment selection (outer) and cut point determination (inner). The cut point uses the Shapley interaction index (Lundberg et al., 2018). The outer optimization enumerates segments to select the one with the smallest interaction index. A segment $w_{(s_{i},s_{i+1})}$’s contribution ($s_{i},s_{i+1}$ are the start and end positions) measures the prediction’s distance from the prediction boundary (label $y$’s contribution). Figure 1 (from (Chen et al., 2020)) contrasts HEDGE and ACD. HEDGE initially splits sentences into larger chunks (e.g., “it never” and “fails to engage us”), while ACD merges words (e.g., “fails” and “to”). In this example, HEDGE’s explanations show that the LSTM makes an incorrect prediction by missing the “never” and “fails” interaction, which ACD fails to capture.
We summarize the approaches discussed in this section, along with the different datasets, and their respective codebase links, in Table 6. In Table 10, the relevant datasets are listed with the specific task, type of description, and dataset statistics.
Approach Explanation type Dataset(s) Codebase
Murdoch and Szlam (2017) word-level interaction of LSTM WikiMovies, Yelp, SST No URL provided
Murdoch et al. (2018) phrase-level interaction of LSTM SST, Yelp github.com/jamie-murdoch/ContextualDecomposition
ACD(Singh et al., 2019) Hierarchical interaction bottom up SST github.com/csinva/hierarchical-dnn-interpretations
Chen et al. (2020) Hierarchical interaction top down SST, IMDB github.com/UVa-NLP/HEDGE
Table 6. Summary table of the explanations for sequence models described in Section 8.
9. Attention
The attention mechanism, originally termed as an alignment model in the context of machine translation by  Bahdanau et al. (2015), has been a key component in text processing using deep neural networks since it was proposed. The attention models used in the literature are of two types: additive attention (Bahdanau et al., 2015) and scalar dot product (Vaswani et al., 2017). We assume a generic model framework and notations (Jain and Wallace, 2019; Mohankumar et al., 2020).
Suppose, we have a sequence of tokens $\{{w}_{1},{w}_{2},\ldots,{w}_{T}\}$. The model inputs $x$ $\in$ $\mathbb{R}^{T\times|V|}$, containing one-hot vectors at each position of the words and $V$ is the vocabulary size. This sequence is passed over an embedding layer to get a $d$ dimensional representations of each tokens $x_{e}$ $\in$ $\mathbb{R}^{T\times d}$. Encoder module takes the representations and produces TTmm dimensional hidden states. Note that one can consider LSTM, RNN or any sequence modeling network as a choice for the encoder module. Formally, the encoded representation is $h=Enc(x_{e})$ $\in$ $\mathbb{R}^{T\times m}$. In a similar way we can encode a query $Q$ $\in$ $\mathbb{R}^{m}$ (e.g., hidden representation of a question in question answering task). A similarity function $\phi$ takes $h$, $Q$ and produces scalar scores. Attention scores can be computed as: $\hat{\alpha}$ = softmax($\phi$($h,Q$)), where $\hat{\alpha} \in$ $\mathbb{R}^{T}$. Context vectors are obtained by combining attention scores with the hidden states. There are mainly two common similarity functions i.e., the choice of $\phi$, i) additive $\phi$($h,Q$) = $v^{T}tanh(W_{1}h+W_{2}Q)$ and ii) scaled dot product $\phi$($h,Q$) = $\frac{hQ}{\sqrt{m}}$, where $W_{1}$, $W_{2}$ are learnable parameters and tanhtanh is the activation function.
Attention distributions can be considered as inherently explainable if there is a cause-effect relationship between higher attention and the model prediction. A higher weight implies a greater impact on the model’s prediction. In general, attention based explanations are characterised as faithful and plausible (Mohankumar et al., 2020). If higher attention signifies a greater impact on the model prediction, it is called as faithful explanation. To evaluate a faithful explanation, we need to understand the correlation between inputs and outputs. One popular evaluation technique is to order the words based on the attention weights and compute the average number of tokens required to flip the decision of a prediction (Nguyen, 2018). On the other hand, the explanation is considered as plausible if its weights can be understood by human. We need humans to evaluate whether they are plausible explanation or not.
9.1. Debate on whether attention is explanation or not
While attention was initially considered inherently explainable, Jain and Wallace (2019) argued against this, conducting experiments across various datasets (SST, IMDB, ADR Tweets, 20 Newsgroups, AG News, Diabetes, Anemia, CNN, bAbI, and SNLI) using both additive and scalar dot-product attention. They found that the correlation between attentions and feature attribution scores (gradient based measure) are weak for BiLSTMs with the attention components. Furthermore, they demonstrated that adversarial attention distributions (permutations of original attention weights) yielded similar results, suggesting attention weights are not reliable as explanations.
Further, Wiegreffe and Pinter (2019) showed that attention weights frozen to a uniform distribution performed comparably to original weights on some datasets, suggesting attention is not essential for these tasks. They showed that Replacing the LSTM with a multi-layered perceptron (MLP) while preserving learned attention weights resulted in significantly better performance than just a trained MLP, indicating that attention captures word-level architecture. An attempt to train an adversarial network with divergent attention distributions proved difficult and task-dependent. All experiments used binary classification tasks with LSTMs on datasets including SST, IMDB, 20 Newsgroups, AG News, Diabetes, and Anemia.
9.2. Does sparse attention improve interpretability?
Recent work has explored sparse attention (Martins and Astudillo, 2016; Malaviya et al., 2018) to reduce computational cost, with the idea that attending to fewer tokens could also improve interpretability. However, Meister et al. (2021) showed that sparseness does not guarantee better interpretability. Using the Sparsegen (Laha et al., 2018) projection function to create sparse attention distributions (controlled by sparsity parameters), they investigated the relationship between input tokens and their co-indexed intermediate representations - more specifically, how much each input token influences the magnitude of these representations. They also checked if these representations could be primarily linked to a single input part, proposing a normalized entropy of gradient-based feature importance scores. A low entropy (e.g., 0) indicates a unique influence. Experiments on IMDB, SST, and 20News (topic classification) using LSTM and Transformer models suggested that the relationship between intermediate representations and input tokens is not one-to-one, even with sparse attention. They found that adding sparsity decreases plausibility and that sparse attention correlates less with other feature importance measures compared to standard attention.
9.3. Faithful and plausible attention distributions by changing the loss function
A body of work has attempted to foster more faithful and plausible attention distributions. One approach incorporates faithfulness by modifying the model’s objective/loss function with a penalty term, applicable to attention distributions and sequence model hidden states. Mohankumar et al. (2020) proposed Orthogonal and Diversity LSTMs, observing that vanilla LSTMs’ hidden states often exhibit similarity, occupying a narrow cone in latent space and yielding similar predictions even with permuted attention weights. Orthogonal LSTM orthogonalizes hidden states using the Gram–Schmidt process, while Diversity LSTM adds a conicity measure as a reward factor to the loss function, penalizing similar hidden state vectors. The conicity measure for a set of vectors $V=\{{v}_{1},{v}_{2},\ldots,{v}_{n}\}$ is defined as:
$\mathit{conicity}(V)=\frac{1}{n}\sum_{i=1}^{n}\cos(v_{i},\frac{1}{n}\sum_{j=1}^{n}v_{j}).$
High conicity indicates low hidden state vector diversity. Evaluating these modified LSTMs on twelve datasets (text classification, NLI, paragraph detection, and QA) showed that Orthogonal and Diversity LSTMs had much lower conicity than vanilla LSTMs. Diversity LSTM’s extracted rationales received more focused attention and were shorter. This model is more human-understandable and correlates well with integrated gradients (Sundararajan et al., 2017), reducing attention on punctuation and improving distribution across parts of speech. Diversity LSTMs focus more on nouns and adjectives, which are important for tasks like sentiment analysis.
Chrysostomou and Aletras modified the Transformer loss function to incorporate faithfulness. Using TEXTRANK (Mihalcea and Tarau, ) to compute word salience distributions, they penalized the model when attention deviated from these distributions using KL divergence. Experiments on SST, AGNews, EV.INF, M.RC, and Semeval showed improved faithfulness compared to rationale extraction and input erasure.
Table 7 provides summary, datasets, and the codebase of the explanation approaches discussed in this section. The details of the datasets related to the attention based explanations can be found in Table 10.
Approach Explanation type Datasets Codebase
Jain and Wallace (2019) Claims attention is not explainable SST, IMDB, ADR Tweets, 20 Newsgroups, AG News, Diabetes, Anemia, CNN, bAbI, SNLI github.com/successar/AttentionExplanation
Wiegreffe and Pinter (2019) Argues attention maybe explainable SST, IMDB, 20 Newsgroups, AG News, Diabetes, Anemia github.com/sarahwie/attention
Meister et al. (2021) Sparsify attention / attend to fewer tokens IMDB, SST, 20 Newsgroups No URL provided
Mohankumar et al. (2020) Faithful and plausible attention with modified LSTM SST, IMDB, Yelp, Amazon, Anemia, Diabetes, 20 Newsgroups, Tweets, SNLI, QQP, babI, CNN github.com/akashkm99/Interpretable-Attention
Chrysostomou and Aletras Faithful transformer based predictions SST, AG News, EV.INF, MultiRC github.com/GChrysostomou/saloss
Table 7. Summary table of the explanations for the attention module discussed in Section 9.
10. Transformers and BERT
Since the invention 66todo: 6invention: too strong a word? of the Transformer (Vaswani et al., 2017) and subsequently BERT (Devlin et al., 2019), the landscape of neural NLP and IR has changed significantly, as BERT based models have been very effective on various NLP and IR tasks. Consequently, a number of researchers have focused on the interpretability of different components of Transformer and BERT. The survey by Rogers et al. (2020) provides an aggregated view of the research done in this topic. Broadly, research questions regarding the explainability of BERT can be categorised into the following classes:  (1) The significance of different layers of BERT,  (2) Investigation on BERT’s capability (or inability) to learn linguistic knowledge,  (3) The role of different attention heads, and  (4) Analyses of the information flow in BERT. In this section, we provide an overview of the research efforts categorized into these aspects. Additionally, we summarize the findings by these works in Table  8.
Findings Citations (Codebases)
BERT-base beats BERT-large on subject-verb agreement and sentence subject detection. Goldberg (2019) (github.com/yoavg/bert-syntax) and  Lin et al. (2019) (github.com/yongjie-lin/bert-opensesame)
BERT struggles to adjust over negation. Ettinger (2020) (github.com/aetting/lm-diagnostics)
BERT does not comprehend arguments. Niven and Kao (2019) (github.com/IKMLab/arct2)
Most of the effectiveness of Transformer and BERT are due to a few of the attention heads, as pruning the rest does not significantly reduce the performence of the models. Voita et al. (2019) (github.com/lena-voita/the-story-of-heads) and Budhraja et al. (2021) (NA)
Attention heads in the earlier layers are local, i.e., tokens attend to nearby tokens. There are some evidences that the attention heads perform different roles to learn different linguistic knowledge. Clark et al. (2019) (github.com/clarkkev/attention-analysis), Pande et al. (2021) (github.com/iitmnlp/heads-hypothesis)
The middle layers of BERT are multi-skilled heads, later layers are specific to task. Pande et al. (2021)
Table 8. A summary of what BERT can do and why, as demonstrated by several research works in Section 10.
10.1. Significance of different layers of BERT
As expected, larger BERT variants (with more layers and hidden states) generally outperform base variants (Devlin et al., 2019), though exceptions exist, such as subject-verb agreement (Goldberg, 2019) and sentence subject detection (Lin et al., 2019), where the base model performs better. These anomalies are rare, and it’s unclear if the larger model is inherently less potent or if further fine-tuning could improve its performance.
Several works investigate what aspects different layers of deep architectures encode, typically using layer-specific word representations in probing tasks to determine captured syntactic representations. Blevins et al. (2018) demonstrated that RNNs encode some syntax without explicit supervision, using probing classifiers on different RNN layers to predict parts-of-speech and constituent label ancestry (parent, grandparent, great-grandparent). Using English Universal Dependencies, CoNLL-2012, WMT02014 English-German, and CoNLL-2012 datasets for dependency parsing, semantic role labeling, machine translation, and language modeling, they trained deep RNNs and observed that higher layers encode less syntactic structure and more abstract representations, based on layer-wise prediction accuracy on various syntactic tasks.
In a spirit similar to (Blevins et al., 2018), Tenney et al. (2019a) proposed two scoring functions to evaluate BERT layers. The first function, namely scalar mixing weight, is designed to evaluate the importance of individual layers. With $l$ layers and $j^{th}$ layer representation $y_{j}$, the final decision is based on $\sum\alpha_{j}y_{j}$, where $\alpha_{j}$ is tuned on training data to determine the importance of layers. Another scoring formula, namely cumulative score, denotes how much the accuracy measure F1 score changes for the introduction of a specific layer of BERT. Term $Score(y_{j})-Score(y_{j-1})$ can be used to measure the changes in F1 score observed while adding the $j^{th}$ layer.
Experiments on English Web Treebank, SPR1, SemEval, and OntoNotes 5.0 for eight tasks (POS, constituents, dependencies, entities, semantic role labeling, coreference, semantic proto-roles, and relation classification) showed BERT follows a similar trend to RNNs: basic syntax is captured earlier, and later layers are better for semantics. Syntactic structures are generally more localizable, while semantics are spread across layers. However, Tenney et al. (2019a) noted cases where syntactic identification occurs in later layers. For example, in “china today blacked out a cnn interview that was…”, initial layers tag “today” as a common noun (date). Due to ambiguity, the model later updates “china today” to a proper noun, changing its semantic role. In this case, the model performs syntactic, then semantic, and finally both syntactic and semantic tasks.
10.2. BERT’s capability (or inability) to learn linguistic knowledge
Jawahar et al. (2019) identified linguistic structures captured by BERT base, confirming that surface structures are encoded in initial layers, syntactic structures in middle layers, and semantic information in upper layers, using probing tasks (SentEval (Conneau and Kiela, 2018)), as discussed in Section 4.1. Experiments across ten probing tasks considered surface features (sentence length), syntactic features (tree depth), and semantic features (tense) (Conneau et al., 2018). BERT also captures dependency structures, with higher layers used for long-range dependencies. McCoy et al. (2019) used a Tensor Product Decomposition Network (TPDN) to validate whether BERT implicitly represents role-filler pairs (e.g., in sequence copying, $\{3,8,7,5\}$ as $3:first+8:second+7:third+5:fourth$). They used syntactic tree position (path from root) as the role and the word as the filler, assuming that if a TPDN can approximate a neural network’s learned representation for a role scheme (by minimizing mean squared error), the network implicitly captures that scheme. Several researchers have investigated BERT’s limitations, showing it struggles with argument comprehension (Niven and Kao, 2019) and negation (Ettinger, 2020).
10.3. Understanding multi-head attention
Vaswani et al. (Vaswani et al., 2017) claimed that multi-head attention allows attending to information from different representation subspaces, unlike single-head attention. Subsequent works investigating Transformer and BERT attention maps have partially confirmed this. Voita et al. (2019) analyzed Transformer attention maps for machine translation, showing that only a few heads are responsible for performance, suggesting many can be pruned. Budhraja et al. (2021) adopted this for multilingual BERT. Clark et al. (2019) investigated BERT-base attention maps, probing each head for specific syntactic relations on dependency parsing datasets. They observed that earlier layers contain heads attending heavily to previous/next tokens. Attention to [SEP] is largely inconsequential, while broad attention occurs in lower layers and the last layer’s [CLS] token. Attention maps represent English syntax, distributed across multiple heads. Some heads behave similarly, and heads within the same layer often have similar distributions. While probing showed some heads focus on specific dependency relations (e.g., coreference, preposition/object, verb/direct object), individual heads do not clearly attend to the overall dependency structure.
In a follow up work, Pande et al. (2021) formalized the analysis by classifying the attention heads based on their functional roles, into the following categories:  (a) local: tokens are attended in a small neighborhood of current token, generally next/previous tokens.  (b) syntactic: attending tokens which are syntactically related to the current token.  (c) delimiter: attending special tokens, i.e., [CLS] and [SEP].  (d) block: tokens are attended within the same sentence contrary to the tokens in the sentences before or after the [SEP] token. Figure 2 shows this categorization with a sample sentence. The authors defined attention sieve as a set of tokens that will be attended to obtain the representation for a specific attention role. They also defined a sieve bias score as the ratio between the average attention scores given to the tokens in the sieve and to all tokens in the input sequence. In case of syntactic attention heads, the set of tokens from the input sequence that make up a sieve was identified by a dependency parser. For other heads it is quite straightforward to determine the sieve tokens. To identify attention heads that satisfy specific functional roles, Pande et al. used thresholds on the sieve bias score. Their proposed framework provides an option for tuning the strictness of the classification of attention heads by adjusting the thresholds. The higher the threshold value, the stricter the classification rule for designating the particular head. Experiments on QNLI (Wang et al., 2018) (QA Natural Language Inference), QQP (Wang et al., 2018) (paraphrase detection), MRPC (Dolan and Brockett, 2005) (paraphrase detection), and SST-2 (Socher et al., 2013) (sentiment analysis)suggest that middle layers of BERT have multi-skilled heads, while later layers are task-specific. Syntactic heads are also local heads and delimiter heads overlap with other functional heads.
Figure 2. An example (reproduced from (Pande et al., 2021)) showing the attention sieve received by various input tokens. The current tokens are marked with boxes and color coded edges point to other tokens in the sieve for different functional roles: a) block, b) nsubj (syntactic), c) local, d) amod (syntactic), and e) delimiter.
10.4. Information flow in BERT
Abnar and Zuidema (2020) attempted to quantify attention flow in Transformers. In a later work, Lu et al. (2021) observed that most of the information is passing through the skip connections of BERT, and showed that attention weights alone are not sufficient enough to quantify information flow. Using a gradient-based feature attribution (Sundararajan et al., 2017) method that has been discussed earlier, Lu et al. (2021) introduced influence patterns as how gradient flows from input to output through skip connections, dense layers, etc. Gradient based influence patterns from input to output token can be constructed considering all possible paths using the chain rule. To reduce the computational overhead, they have used strategies to greedily add influential nodes from one layer to the adjacent layer. A path from input tokens to the output is considered a form of explanation. Experimental findings were conducted on subject word agreement and reflexive anaphora tasks. While previous methods only look into the aspects of attention weights, they do not discuss how information is traveling through the skip connections. In this work, they argued that BERT uses attention to contextualize information and skip connections to pass that information from one layer to another. Once contextualize step is finalized it relies more upon the skip connections. Experimental results also show that BERT is able to do negation in sentiment analysis tasks very well.
Another information interaction in the Transformer network was studied by Hao et al. (2021) through the lens of self-attention attribution (Sundararajan et al., 2017) based method. With feature attribution (integrated gradients) techniques, they identified important attention connections i.e. each pair of words in the attention heads of different layers. Intuitively, pairs of words receive a high score if they contribute more to the final prediction. It was observed that a larger attention score between a pair of words may not contribute towards the final prediction.
With the attribution scores of different heads, they pruned attention heads from different layers of BERT. Accuracy values do not change much if the heads with small attribution scores were pruned; however, heads with large attribution scores significantly impact the performance. In general, it was observed that two heads across each layer produce a good performance, and pruning the heads with average attention scores does not significantly affect performance. Further, important attention heads are highly correlated across different datasets. All the empirical results were conducted on MNLI, RTE, SST-2, and MRPC datasets. A heuristic algorithm was designed to demonstrate the information flow between input tokens inside the BERT layer. Figure 3 shows two instances, one from the MNLI dataset and another from the sentiment analysis dataset SST-2. For the first one, the interactions between input sentences try to explain the prediction. However, for the second example, although ‘seldom’, ‘movie’, and ‘man’ are more important words for the ‘positive’ sentiment to a human; the information flows presented by their approach do not aggregate to these tokens.
(a) Attribution tree generated from MNLI dataset, where the final prediction is ‘entailment’.
(b) Attribution tree generated from SNLI-2 dataset; the prediction of the sentence is ‘positive’.
Figure 3. Reproducing figures from Hao et al. (2021), showing the examples of attribution trees generated with different datasets.
11. Future Research
In the preceding sections, we have identified several issues and research questions that have not been adequately addressed so far. We conclude in this section by listing some of these open questions.
First, it is important to investigate how the nature of NLP tasks differs from that of more traditional IR tasks. Traditional IR datasets (e.g., Robust04, ClueWeb09b) and QA-oriented datasets (e.g., MSMARCO) are clearly different with respect to document sizes, query nature (keyword vs. QA), and the presence/absence of $\langle Q,D_{+},D_{-}\rangle$ triplets (where $D_{+}$ and $D_{-}$ are positive and negative relevant documents for a query $Q$). These differences, combined with the fact that the Transformer models at the core of NRMs are trained for NLP tasks, raise questions about whether the notable success of NRMs is confined to QA-oriented datasets. Indeed, several pre-BERT NRMs have been shown to struggle on datasets such as Robust04 and LETOR4.0 (Pang et al., 2017). In addition, BERT-based models have not been adequately tested on Web collections, e.g., ClueWeb09b and Gov2, because of their scale. One reason behind the limitation of BERT-based retrieval models may be tokenization, as standard tokenizers may not be suitable for all corpora. For example, TREC-DL 2019’s medical queries necessitate adaptive tokenizers, as explored by Balde et al. (2024) using adaptive byte-pair encoding. In general, a more comprehensive failure analysis of BERT-based models is needed. On a related note, domain-specific explainability in areas like medical and legal retrieval presents unique challenges requiring further research (Polley, 2022).
As discussed in Section 3.2, Moshkovitz et al. (2020) highlighted the difficulty of explaining cluster assignments in high-dimensional spaces, proposing visualization using axis-aligned features and shallow threshold trees. Applying this to explain rankings from NRMs using decision tree-based explainers is a promising direction. Given that NRMs are optimized by minimizing a loss function over $(Q,D_{+},D_{-})$ triplets, it is worth investigating whether we can quantify the influence of specific triplets on the loss function. This could reduce training costs, improve explanations, and result in effective mining of negative documents.
At a more basic level, the interpretability of IR evaluation measures also warrants investigation. While some research addresses ranking constraints for metric axiomatics [^7] https://www.eecis.udel.edu/~hfang/AX.html, the interpretability of metrics such as Average Precision and NDCG remains under-explored, especially given NDCG’s non-linear nature.
Arguably the most pressing need of the hour, however, is a widely accepted evaluation framework for Explainable IR/NLP, consisting of standardised experimental protocols and test collections (see Section 5). This would enable reliable and reproducible comparisons of different explanation methods and foster further advances in the field. Given the increasing emphasis on understanding complex models, and emerging regulatory requirements, we hope that standard IR evaluation forums (e.g., TREC, CLEF and NTCIR) will work towards fulfilling this requirement for different categories of explanation methods.
Appendix A Supplementary Tables (electronic version only)
Dataset Collection Type #Docs #Queries Pool Depth
MSMARCO passage (train) Question answering 8,841,823 502,939 Shallow
MSMARCO passage (dev) 6,980 Shallow
MSMARCO passage (test) 6,837 Shallow
MSMARCO document (train) Web 3,213,835 367,013 Shallow
MSMARCO document (dev) 5,193 Shallow
MSMARCO document (test) 5,793 Shallow
TREC DL passage 2019 [Craswell et al., 2020] Question answering 8,841,823 43 Deep
TREC DL document 2019 [Craswell et al., 2020] Web 3,213,835 43 Deep
TREC DL passage 2020 [Craswell et al., 2021] Question answering 8,841,823 54 Deep
TREC DL document 2020 [Craswell et al., 2021] Web 3,213,835 45 Deep
Robust News 528,155 249 Deep
Clueweb09b Web 50,220,423 200 Deep
Table 9. Some statistics about datasets commonly used for experiments related to the interpretability of document ranking tasks.
Dataset Descriptions Task Statistics
SST Movie Reviews Sentiment analysis 6,920 / 872 / 1,821
IMDB Movie Reviews Sentiment analysis 25,000 / 4,356
WikiMovies Movie Domain Question Answering 100,000
Yelp Review Polarity Business Reviews Sentiment analysis 560,000 / 38,000
M.RC Reading Comprehension Question Answering 24,029 / 3,214 / 4,848
SEMEVAL Tweets Classification 6,000 / 2,000 / 20,630
Ev.Inf Biomedical Articles Inference 5,789 / 684 / 720
AG News Classification 102,000 / 18,000 / 7,600
QQP Quora Questions Paraphrase 364,000 / 391,000
QNLI Wikipedia QA / Natural Language Inference 105,000 / 5,400
MRPC News Paraphrase 3,700 / 1,700
ADR Tweets Twitter Adverse Drug Reaction Classification 16,385 / 4,123
20 Newsgroups News Classification 1,426 / 334
AG News News Classification 60,000 / 3,800
Diabetes (MIMIC) Clinical Health Classification 7,734 / 1,614
Anemia (MIMIC) Clinical Health Classification 5,098 / 1,262
CNN News Question Answering 380,298 / 3,198
bAbI (Task 1 / 2 / 3) Supporting Facts Question Answering 10,000 / 1,000
SNLI Human-written Sentence Pairs Natural Language Inference 549,367 / 9,824
Table 10. Details of the datasets mentioned in this survey for explanations of sequence models (§ 8) and attention (§ 9).